OpenStudy (anonymous):

solve for t, -4(cos2t)(sin2t) = 0

OpenStudy (anonymous):

Can't you use the sin2x identity?

OpenStudy (anonymous):

probably, but haven't gotten there yet, soo hahah

sam (.sam.):

-4cos2t+sin2t=0 -4(1-2sin^2t)(2sintcost)=0 -4(2sintcost-4sin^3tcost)=0 -8sintcost+16sin^3tcost=0 2sin^3tcost=sintcost sin^2t=1/2 sint=sqrt2/2 t=45,135

OpenStudy (anonymous):

the answer is npi/4, where n is an integer

sam (.sam.):

thats in radian, 45degrees = pi/4

OpenStudy (anonymous):

if only i could understand your work

sam (.sam.):

where you dont understand?

OpenStudy (anonymous):

Do you have a formula sheet you can use? If you know the Sin(2x) identity, then it would be a lot simpler .

OpenStudy (anonymous):

no, it should just be algebra